calculate the number of mol of helium in a 1.20-L balloon at 29C and 2.10 atm of pressure.
answer in scientific notation.
1
Expert's answer
2015-10-28T03:57:46-0400
The ideal gas law is:
PV=nRT
where the letters denote pressure(P), volume(V), amount (n, in moles), ideal gas constant(R), and temperature of the gas(T).
Then
n = PV/RT = 2.10*1.20/(0.082*(273+29)) = 0.102 moles